OpenStudy (aimmebgood):

3a^(2) - 21a + 30 factor completely.

OpenStudy (anonymous):

3(a^2 - 7a +10)=3(a^2 - 5a -2a +10)=3(a(a-5) -2(a-5))=3(a-5)(a-2)

OpenStudy (campbell_st):

take 3 as a common factor \[3(a^2 - 7a + 10)\] makes it a lot easier

OpenStudy (anonymous):

this time have you understood how it becomes -7a = -5a - 2a @aimmeBgood

OpenStudy (anonymous):

hello \[3(a-2)(a-5)\]